Two people can swing on a trapeze at the same time as long as their total weight is less than 500 pounds. Joe weighs 240 pounds and is 6 feet tall. Jake weight 250 pounds and is 32 years old. Can Joe and Jake swing on the trapeze at the same time?
A. Too little
B. Too much

Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

total weight = 240 + 250 = 490
490 is less than 500
---
answer
yes, they can swing at the same time
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by stanbon(75887)

Two people can swing on a trapeze at the same time as long as their total weight is less than 500 pounds. Joe weighs 240 pounds and is 6 feet tall. Jake weight 250 pounds and is 32 years old.
Can Joe and Jake swing on the trapeze at the same time?
240+250 = 490
Ans: yes
-------------------
Cheers,
Stan h.
------------------